OpenStudy (anonymous):

A local movie theatre sells popcorn in a rectangular prism box.The Smaller box can hold 192 cubic inches of popcorn.As a special promotion,the movie theatre decides to double all the dimensions of the box and sell the popcorn for for the same price as the small box.How much popcorn will the new box hold?

OpenStudy (bahrom7893):

Volume of a rectangular prism: V = l * w * h (length times width times height) V = 192 => l*w*h = 192.

OpenStudy (anonymous):

thx

OpenStudy (bahrom7893):

If you double every dimension, then: 2L * 2W * 2H = V.

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

Which means that your total volume will increase by 6

OpenStudy (anonymous):

you beat me to it bahroom

OpenStudy (bahrom7893):

So new amount to be held is: 6 * 192 = 1152

OpenStudy (anonymous):

Hey can you hell me with all my hw today